# MIVE: New Design and Benchmark for Multi-Instance Video Editing

### v1

@@ C.3 Local Metrics Computation
D Comparison with State-of-the-Art Methods
D.1 Qualitative Comparison
Figure 12: Qualitative comparison with SOTA video editing methods (Part 1). (a) shows the color-coded masks overlaid on the input frames to match the corresponding instance captions. (b)-(f) use global target captions for editing. (g) uses global and instance target captions along with bounding boxes (omitted in (a) for better visualization). Our MIVE in (h) uses instance captions and masks. Unfaithful editing examples are shown in red arrow and attention leakage are shown in green arrow.
Figure 13: Qualitative comparison with SOTA video editing methods (Part 2). (a) shows the color-coded masks overlaid on the input frames to match the corresponding instance captions. (b)-(f) use global target captions for editing. (g) uses global and instance target captions along with bounding boxes (omitted in (a) for better visualization). Our MIVE in (h) uses instance captions and masks. Unfaithful editing examples are shown in red arrow and attention leakage are shown in green arrow.
Figure 14: Qualitative comparison with SOTA video editing methods (Part 3). (a) shows the color-coded masks overlaid on the input frames to match the corresponding instance captions. (b)-(f) use global target captions for editing. (g) uses global and instance target captions along with bounding boxes (omitted in (a) for better visualization). Our MIVE in (h) uses instance captions and masks. Unfaithful editing examples are shown in red arrow and attention leakage are shown in green arrow.
We present more visual comparison in Figures 12, 13, and 14. The videos are selected with increasing difficulty based on the number of instances to edit.
D.2 Quantitative Results Based on Instance Sizes and Numbers of Instances
@@
Scope LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA $\uparrow$
ControlVideo [70] ICLR’24 Global 0.9546 0.1684 0.3479 0.3875 0.9516 0.1852 0.3248 0.5220 0.9580 0.2048 0.5845 0.5003
FLATTEN [11] ICLR’24 Global 0.9519 0.1789 0.4230 0.4215 0.9457 0.1902 0.2338 0.5766 0.9547 0.1876 0.2371 0.4949
RAVE [26] CVPR’24 Global 0.9547 0.1752 0.3953 0.4310 0.9527 0.1830 0.2932 0.5444 0.9581 0.1910 0.3684 0.4815
TokenFlow [16] ICLR’24 Global 0.9486 0.1783 0.4517 0.4434 0.9406 0.1876 0.3297 0.6085 0.9522 0.1880 0.3536 0.5090
FreSCo [62] CVPR’24 Global 0.9288 0.1790 0.4044 0.4283 0.9226 0.1892 0.2937 0.5945 0.9383 0.1852 0.2709 0.4958
GAV [25] ICLR’24 Local, Multiple 0.9529 0.1803 0.4224 0.4680 0.9498 0.1932 0.3298 0.5913 0.9550 0.1889 0.3740 0.5420
MIVE (Ours) - Local, Multiple 0.9537 0.1794 0.4051 0.6059 0.9441 0.1997 0.4647 0.6883 0.9509 0.2243 0.7414 0.7331
@@
Scope GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ BP $\downarrow$
Editing on 1-3 Instances (Easy Video) - 116 Videos
ControlVideo [70] ICLR’24 Global 0.9729 0.2724 0.8839 0.9513 0.2020 0.5374 0.6192 79.3548
FLATTEN [11] ICLR’24 Global 0.9661 0.2415 0.2924 0.9484 0.1893 0.2568 0.6060 67.1713
RAVE [26] CVPR’24 Global 0.9661 0.2698 0.5452 0.9533 0.1886 0.3620 0.5964 71.2598
TokenFlow [16] ICLR’24 Global 0.9686 0.2578 0.5710 0.9465 0.1880 0.3480 0.6248 74.2420
FreSCo [62] CVPR’24 Global 0.9534 0.2491 0.4138 0.9327 0.1877 0.2841 0.6085 92.0291
GAV [25] ICLR’24 Local, Multiple 0.9643 0.2518 0.5583 0.9477 0.1915 0.3851 0.6466 63.3588
MIVE (Ours) - Local, Multiple 0.9583 0.2738 0.8589 0.9441 0.2203 0.6932 0.7983 58.4894
Editing on 4-7 Instances (Medium Video) - 66 Videos
ControlVideo [70] ICLR’24 Global 0.9757 0.2775 0.8834 0.9579 0.1875 0.4543 0.3704 65.3667
FLATTEN [11] ICLR’24 Global 0.9700 0.2373 0.2305 0.9526 0.1853 0.2288 0.4199 56.7246
RAVE [26] CVPR’24 Global 0.9689 0.2777 0.5819 0.9570 0.1842 0.3396 0.3951 57.2210
TokenFlow [16] ICLR’24 Global 0.9686 0.2559 0.5425 0.9486 0.1845 0.3631 0.4498 62.3709
FreSCo [62] CVPR’24 Global 0.9555 0.2589 0.4276 0.9318 0.1835 0.3177 0.4389 76.8818
GAV [25] ICLR’24 Local, Multiple 0.9674 0.2648 0.5549 0.9549 0.1859 0.3660 0.4676 57.8612
MIVE (Ours) - Local, Multiple 0.9614 0.2763 0.8397 0.9501 0.2060 0.5872 0.6483 48.2502
Editing on $>$7 Instances (Hard Video) - 18 Videos
ControlVideo [70] ICLR’24 Global 0.9781 0.2692 0.9051 0.9651 0.1885 0.3602 0.1698 58.5803
FLATTEN [11] ICLR’24 Global 0.9717 0.2274 0.2007 0.9584 0.1898 0.2499 0.2339 57.0569
RAVE [26] CVPR’24 Global 0.9711 0.2735 0.7714 0.9602 0.1857 0.3242 0.2026 51.7408
TokenFlow [16] ICLR’24 Global 0.9682 0.2552 0.5778 0.9529 0.1878 0.3162 0.2212 55.8445
FreSCo [62] CVPR’24 Global 0.9538 0.2536 0.4339 0.9322 0.1844 0.2954 0.2156 71.5167
GAV [25] ICLR’24 Local, Multiple 0.9715 0.2580 0.4825 0.9628 0.1870 0.2911 0.2210 47.0554
MIVE (Ours) - Local, Multiple 0.9697 0.2784 0.8937 0.9626 0.2002 0.5118 0.3669 49.9864
@@
Table 6:  Quantitative comparison for multi-instance video editing on various number of instances. We categorize 200 videos of MIVE Dataset depending on the number of edited instances: (i) Easy Video (EV): video that contains 1-3 edited instances, (ii) Medium Video (MV): video that contains 4-7 edited instances, and (iii) Hard Video (HV): video that contains $>$ 7 edited instances. The best and second best scores are shown in red and blue, respectively.
@@
In this part, we analyze whether the baseline methods and our MIVE exhibit biases based on: (i) the instance sizes or (ii) the number of instances.
@@
As shown in Tab. 5, the Local Temporal Consistency scores align with the findings in the main paper (Sec. 5.1), where RAVE and ControlVideo achieve the best and second-best performance, respectively, while our MIVE method demonstrates competitive results.
For medium and large instance sizes, our method achieves the best performance in textual faithfulness (LTF and IA) and attention leakage (CIA). In the medium-sized instance scenario, our MIVE shows LTF, IA, and CIA improvement of approximately 0.006, 13.5%, and 8%, respectively, compared to the second-best method. Similarly, in the large-sized instance scenario, MIVE achieves an improvement of approximately 0.02, 16%, and 19%, respectively, over the second-best method. The greater improvement in large-sized instances indicates that our method performs better on larger objects likely due to our reliance on masks as a condition for editing. This reasoning is further supported by the results of GAV. GAV, which uses bounding boxes, tends to have more overlaps in large-instance scenarios, leading to worse performance compared to the global editing baselines. These findings highlight the advantage of using masks over bounding boxes as conditions for multi-instance video editing.
In the small instance size scenario, the LTF scores of all methods tend to be lower compared to medium and large instances. This highlights the challenges of editing small instances in diffusion-based video editing methods, which are caused by the downsampling in the VAE of the LDM [44]. For small instances, our MIVE achieves the best performance in attention leakage (CIA) and competitive performance in editing faithfulness (LTF and IA). For LTF, the difference between our method and the best-performing method (GAV) is minimal (0.001). It is worth noting that the LTF of our method for small instances could potentially be improved by increasing the value of $\lambda$ in our Instance-centric Probability Redistribution (IPR). However, for IA, our MIVE scores 5% lower than the best method, despite achieving higher LTF. To better understand this issue, we analyze the qualitative results of TokenFlow, as shown in our demo and in Fig. 12, Fig. 13, and Fig. 14. TokenFlow often produces artifacts in multi-instance editing scenarios. For example, in video 2 (middle column) of Fig. 13, where the target instance caption is “baseball” (source instance caption: golf ball), TokenFlow transforms the golf ball into a grassy area, while other methods fail to edit this region, thereby preserving the original golf ball. Interestingly, the IA score of TokenFlow for this edited golf ball is 0.87, whereas our MIVE scores 0.5 even though our results look more like baseball. This observation suggests that the IA metric may not be well-suited for evaluating performance in the scenarios with small-sized instances.
Quantitative results based on number of instances. In addition to the quantitative comparisons with respect to various instance sizes, we provide further analysis for the different numbers of edited instances in each video, as shown in Tab. 6. Videos are categorized into three groups: (i) easy video (EV) having 1–3 edited instances, (ii) medium video (MV) having 4–7 edited instances, and (iii) hard video (HV) having 7 or more edited instances. For the Leakage Scores, there is no significant deviation from the quantitative comparisons in the main paper (Tab. 2). In most cases, our MIVE achieves the best Leakage Scores performance, while either GAV or FLATTEN secures the second-best performance. This aligns with the preferences highlighted in our User Study from the main paper. For Local Scores, the results are consistent with the main paper: our MIVE achieves the best LTF and IA scores, as well as competitive LTC performance. Similarly, for Global Scores, the GTC and FA metrics are also in line with the main paper, with our method achieving competitive GTC and second-best FA results. One notable deviation from the main paper is in the GTF score. While our MIVE achieves the best GTF score in the main paper, this only holds for the EV and HV scenarios. In the MV scenario, RAVE achieves the best result, followed by ControlVideo, with our MIVE ranking third. The margin between our method and the top-performing method is small (0.0014). This deviation in GTF is likely due to CLIP’s limitations in compositional reasoning [23, 36], which hinder its ability to capture complex relationships between instances in the global caption and frames. This further shows the importance of Local Scores in quantifying the performance in multi-instance video editing.
@@
In summary, while there are minor deviations in some scores compared to the quantitative results in the main paper, the performance of our MIVE remains consistent across various scenarios. This demonstrates the robustness of our approach against variations in instance sizes and the number of edited instances.
D.3 User Study Details

### v2

@@ B.3 Local Metrics Computation
C Comparison with State-of-the-Art Methods
Figure 12: Qualitative comparison for three videos (with increasing difficulty from left to right) in our MIVE dataset. (a) shows the color-coded masks overlaid on the input frames to match the corresponding instance captions. (b)-(f) use global target captions for editing. (g) uses global and instance target captions along with bounding boxes (omitted in (a) for better visualization). (h) uses masks and global and local target captions. Our PRIMEdit in (i) uses instance captions and masks. Unfaithful editing examples are shown in red arrow and attention leakage are shown in green arrow.
C.1 Full Qualitative Comparison and Demo Videos
We present the full qualitative comparison in Fig. 12. Notably, TokenFlow [20] and FreSCo [63] are still susceptible to unfaithful editing as shown in the red arrows in Fig. 12-(f) and (g), respectively.
We provide sample input and edited videos for our method along with baseline methods in our project page: https://kaist-viclab.github.io/primedit-site/.
Scope GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ TC $\uparrow$ TF $\uparrow$ Leakage $\uparrow$
ControlVideo [72] ICLR’24 Global 0.9745 0.2738 0.8850 0.9548 0.1960 0.4944 0.4963 0.5327 0.4242 0.0167 0.0656 0.0208
FLATTEN [14] ICLR’24 Global 0.9678 0.2389 0.2627 0.9507 0.1881 0.2463 0.5109 0.8220 0.1314 0.3260 0.0594 0.1646
RAVE [30] CVPR’24 Global 0.9675 0.2728 0.5769 0.9551 0.1869 0.3504 0.4950 0.7699 0.2188 0.0667 0.0354 0.0458
TokenFlow [20] ICLR’24 Global 0.9686 0.2571 0.5617 0.9477 0.1868 0.3499 0.5319 0.7991 0.1807 0.0125 0.0521 0.0531
FreSCo [63] CVPR’24 Global 0.9543 0.2528 0.4202 0.9324 0.1860 0.2960 0.5172 0.7567 0.2427 0.0823 0.0073 0.0104
GAV [28] ICLR’24 Local, Multiple 0.9665 0.2574 0.5544 0.9518 0.1895 0.3733 0.5516 0.8618 0.0914 0.0792 0.0938 0.1000
VideoGrain [65] ICLR’25 Local, Multiple 0.9600 0.2803 0.7430 0.9423 0.2026 0.4798 0.5868 0.8929 0.0488 0.0802 0.1177 0.1542
PRIMEdit (Ours) - Local, Multiple 0.9677 0.2632 0.7707 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586 0.3365 0.5687 0.4510
Table 5:  Full quantitative comparison for multi-instance video editing. The best and second best scores are shown in red and blue, respectively.
C.2 Full Quantitative Results
We present the full quantitative comparison in Tab. 5. Despite achieving the highest scores in local metrics, leakage minimization, background preservation, and user studies, PRIMEdit exhibits a slight disadvantage in global metrics. Our GTC score ranks fourth, which we attribute to prior SOTA methods favoring frame reconstruction, thereby preserving the high GTC of the original video. Similarly, our GTF ranks fourth, while our FA ranks second only to ControlVideo. We attribute this to CLIP’s limitations in compositional reasoning [27, 38], which hinders its ability to capture complex relationships between instances in the global caption and frames. As a result, CLIP-based evaluations may lead to inflated GTF and FA scores despite inaccuracies in instance edits (e.g., swapped edits between two people in Video 2 of Fig. 12-(b)).
C.3 Quantitative Results Based on Instance Sizes and Numbers of Instances
@@
Scope LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$
ControlVideo [72] ICLR’24 Global 0.9202 0.1610 0.3308 0.9552 0.1993 0.5258 0.9554 0.2008 0.5338
FLATTEN [14] ICLR’24 Global 0.9128 0.1707 0.3228 0.9515 0.1890 0.2366 0.9521 0.1891 0.2285
RAVE [30] CVPR’24 Global 0.9255 0.1710 0.3685 0.9552 0.1898 0.3580 0.9557 0.1898 0.3510
TokenFlow [20] ICLR’24 Global 0.9209 0.1701 0.4015 0.9490 0.1883 0.3451 0.9490 0.1887 0.3405
FreSCo [63] CVPR’24 Global 0.9021 0.1681 0.3686 0.9347 0.1869 0.2948 0.9348 0.1867 0.2825
GAV [28] ICLR’24 Local, Multiple 0.9182 0.1716 0.3358 0.9514 0.1907 0.3810 0.9521 0.1908 0.3807
VideoGrain [65] ICLR’25 Local, Multiple 0.9191 0.1750 0.4419 0.9427 0.2052 0.4991 0.9427 0.2065 0.5025
PRIMEdit (Ours) - Local, Multiple 0.9235 0.1728 0.3870 0.9551 0.2078 0.5502 0.9556 0.2093 0.5604
@@
Scope GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Editing on One (1) Instance (Easy Video) - 200 Videos
ControlVideo [72] ICLR’24 Global 0.9755 0.2717 0.6922 0.9611 0.2048 0.5351 1.0000 0.4130 0.5090
FLATTEN [14] ICLR’24 Global 0.9699 0.2538 0.3140 0.9586 0.1826 0.1971 1.0000 0.7783 0.1567
RAVE [30] CVPR’24 Global 0.9691 0.2774 0.4243 0.9599 0.1857 0.3374 1.0000 0.7044 0.2716
TokenFlow [20] ICLR’24 Global 0.9715 0.2621 0.4413 0.9564 0.1822 0.3006 1.0000 0.7546 0.2148
FreSCo [63] CVPR’24 Global 0.9563 0.2630 0.3200 0.9423 0.1805 0.2515 1.0000 0.6894 0.3014
GAV [28] ICLR’24 Local, Multiple 0.9701 0.2537 0.4453 0.9584 0.1851 0.3350 1.0000 0.8340 0.0999
VideoGrain [65] ICLR’25 Local, Multiple 0.9659 0.2696 0.5474 0.9515 0.2020 0.4531 1.0000 0.8731 0.0570
PRIMEdit (Ours) - Local, Multiple 0.9737 0.2534 0.6192 0.9614 0.2108 0.5562 1.0000 0.8807 0.0661
Editing on 1-3 Instances (Medium Video) - 116 Videos
ControlVideo [72] ICLR’24 Global 0.9730 0.2724 0.8839 0.9512 0.2020 0.5380 0.6191 0.4879 0.4515
FLATTEN [14] ICLR’24 Global 0.9659 0.2417 0.2908 0.9484 0.1894 0.2557 0.6053 0.7902 0.1504
RAVE [30] CVPR’24 Global 0.9661 0.2699 0.5437 0.9532 0.1886 0.3604 0.5971 0.7396 0.2379
TokenFlow [20] ICLR’24 Global 0.9686 0.2580 0.5701 0.9463 0.1880 0.3476 0.6265 0.7720 0.1974
FreSCo [63] CVPR’24 Global 0.9535 0.2492 0.4140 0.9326 0.1877 0.2836 0.6084 0.7165 0.2811
GAV [28] ICLR’24 Local, Multiple 0.9649 0.2528 0.5648 0.9481 0.1919 0.3903 0.6498 0.8397 0.0959
VideoGrain [65] ICLR’25 Local, Multiple 0.9596 0.2775 0.7486 0.9413 0.2065 0.5238 0.6712 0.8736 0.0526
PRIMEdit (Ours) - Local, Multiple 0.9658 0.2629 0.7897 0.9516 0.2096 0.5817 0.7567 0.8823 0.0633
Editing on 4-7 Instances (Hard Video) - 66 Videos
ControlVideo [72] ICLR’24 Global 0.9760 0.2774 0.8815 0.9581 0.1874 0.4544 0.3694 0.5830 0.3920
FLATTEN [14] ICLR’24 Global 0.9702 0.2372 0.2303 0.9527 0.1854 0.2288 0.4206 0.8585 0.1099
RAVE [30] CVPR’24 Global 0.9690 0.2776 0.5821 0.9571 0.1843 0.3399 0.3954 0.8060 0.1957
TokenFlow [20] ICLR’24 Global 0.9686 0.2559 0.5427 0.9487 0.1845 0.3631 0.4504 0.8342 0.1585
FreSCo [63] CVPR’24 Global 0.9557 0.2589 0.4274 0.9319 0.1835 0.3179 0.4391 0.8068 0.1934
GAV [28] ICLR’24 Local, Multiple 0.9679 0.2652 0.5557 0.9553 0.1861 0.3659 0.4692 0.8884 0.0837
VideoGrain [65] ICLR’25 Local, Multiple 0.9603 0.2836 0.6938 0.9429 0.1972 0.4185 0.5096 0.9180 0.0425
PRIMEdit (Ours) - Local, Multiple 0.9691 0.2622 0.7039 0.9578 0.1990 0.4897 0.6087 0.9229 0.0509
Editing on $>$7 Instances (Extreme Video) - 18 Videos
ControlVideo [72] ICLR’24 Global 0.9781 0.2692 0.9051 0.9651 0.1885 0.3602 0.1698 0.6374 0.3660
FLATTEN [14] ICLR’24 Global 0.9717 0.2274 0.2007 0.9584 0.1898 0.2499 0.2339 0.8930 0.0881
RAVE [30] CVPR’24 Global 0.9711 0.2735 0.7714 0.9602 0.1857 0.3242 0.2026 0.8323 0.1806
TokenFlow [20] ICLR’24 Global 0.9682 0.2552 0.5778 0.9529 0.1878 0.3162 0.2212 0.8455 0.1545
FreSCo [63] CVPR’24 Global 0.9538 0.2536 0.4339 0.9322 0.1844 0.2954 0.2156 0.8319 0.1760
GAV [28] ICLR’24 Local, Multiple 0.9715 0.2580 0.4825 0.9628 0.1870 0.2911 0.2210 0.9070 0.0912
VideoGrain [65] ICLR’25 Local, Multiple 0.9615 0.2873 0.9286 0.9478 0.1966 0.4050 0.2516 0.9348 0.0463
PRIMEdit (Ours) - Local, Multiple 0.9750 0.2682 0.8937 0.9689 0.1945 0.4209 0.3413 0.9390 0.0559
@@
Table 7:  Quantitative comparison for multi-instance video editing on various number of instances. We categorize 200 videos of MIVE Dataset depending on the number of edited instances: (i) Easy Video (EV): video that contains 1 edited instance, (ii) Medium Video (MV): video that contains 1-3 edited instances, (iii) Hard Video (HV): video that contains 4-7 edited instances, and (iv) Extreme Video (XV): video that contains $>$ 7 edited instances. The best and second best scores are shown in red and blue, respectively.
@@
In this part, we analyze whether the baseline methods and our PRIMEdit exhibit biases based on: (i) the instance sizes or (ii) the number of instances.
@@
As shown in Tab. 6, the Local Temporal Consistency scores closely align with the findings in the main paper (Sec. 5.1).
For medium and large instance sizes, our method achieves the highest performance in textual faithfulness (LTF and IA) and the second-best in temporal consistency. In these scenarios, both our PRIMEdit and VideoGrain show significant improvements over other SOTA methods in terms of LTF and IA. Notably, in certain cases, global video editing methods outperform GAV. These results underscore the advantage of using masks over bounding boxes as conditioning mechanisms for multi-instance video editing, enabling more precise and context-aware modifications.
In the small instance size scenario, the LTF scores of all methods tend to be lower compared to the medium and large instances. This highlights the challenges of editing small instances in diffusion-based video editing methods, caused by the downsampling in the VAE of the LDM [46].
Quantitative results based on number of instances. In addition to the quantitative comparisons with respect to various instance sizes, we provide further analysis for the different numbers of edited instances in each video, as shown in Tab. 7. Videos are categorized into four groups: (i) easy video (EV) having only one edited instance, (ii) medium video (MV) having 1–3 edited instances, (iii) hard video (HV) having 3-4 edited instances, and (iv) extreme video (XV) having 7 or more edited instances.
For the Leakage Scores, there is no significant deviation from the quantitative comparisons in the main paper (Tab. 2). In most cases, our PRIMEdit achieves the best Leakage Scores performance, while VideoGrain always secures the second-best performance. For Local Scores, the results are consistent with the main paper: our PRIMEdit almost always achieves the best LTC, LTF, and IA scores. Similarly, the global scores align with Tab. 5, where our method achieves competitive GTC and secures the second-best FA performance.
Notably, our PRIMEdit’s performance gain is not solely due to multi-instance editing. PRIMEdit also excels in single-instance editing, as evidenced in the results of EV, where we ourperform baselines specifically optimized for single-object editing across all local metrics. This underscores PRIMEdit’s versatility and effectiveness in delivering localized and faithful edits across various scenarios.
@@
In summary, while there are minor deviations in some scores compared to the quantitative results in Tab. 5 and in Tab. 7, the performance of our PRIMEdit remains consistent across various scenarios. This demonstrates the robustness of our approach against variations in instance sizes and the number of edited instances.
C.4 User Study Details